OpenStudy (anonymous):

-5/2X+1/2=-18 SOLVE pleasee

OpenStudy (anonymous):

first multiply both sides by 2.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

-2x+1=-36

OpenStudy (anonymous):

wait...is -5x

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

-5x+1=-36 -5x=-36-1 -5x=-37 divide both sides by -5 to get x x=-37/5

OpenStudy (anonymous):

cool thanks!

OpenStudy (anonymous):

ur welcome :)